Find the distance, d across the river. Assume the ratio of d to 75 ft is the same as the ratio of 40 ft to 20 ft. The distance across the river is _____ feet. is it 37.50?
Answer #1
It even tells you what the equation is ....

"the ratio of d to 75 ft is the same as the ratio of 40 ft to 20 ft." ---->d/75 = 40/20

d/75 = 2
d = 2(75) = 150
answered by: CheezyReezy